if 2 is added to five times a number, the result is equal to 5 more than four times the number. Find the number.
Answer by checkley77(12844)

5x+2=4x+5
5x-4x=5-2
x=3 ans.
Proof:
5*3+2=4*3+5
15+2=12+5
17=17